OpenStudy (anonymous):

find the value of 10 sqrt 2/ sqrt 5

OpenStudy (lgbasallote):

\[\frac{10\sqrt 2}{\sqrt 5} \implies \frac{10\sqrt 2}{\sqrt 5} \times \frac{\sqrt 5}{\sqrt 5}\] does that help @NYCbaby ?

OpenStudy (jiteshmeghwal9):

Do u gt it @NYCbaby